# This Candidate is [MASK]

### v1

@@ Abstract
I implement a prompt-based learning strategy to extract measures of sentiment and other features from confidential reference letters. I show that the contents of reference letters is clearly reflected in the performance of job market candidates in the Economics academic job market. In contrast, applying traditional “bag-of-words” approaches produces measures of sentiment that, while positively correlated to my LLM-based measure, are not predictive of job market outcomes. Using a random forest, I show that both letter quality and length are predictive of success in the job market. Letters authored by advisers appear to be as important as those written by other referees.
JEL classification:
Key words: Large language models; text data; sentiment analysis.
@@ 1 Introduction
Accessing job seekers’ social networks through references is an important way employers can expedite hiring and reduce search costs. Hiring in real-world labor markets is a costly process characterized by considerable uncertainty regarding the aptitude, skill, and motivation of job applicants. Firms can advertise job vacancies widely and then proceed to a meticulous process of selection through interviewing and testing candidates. Alternatively, firms can economize on search costs by eliciting recommendations from trusted sources. References can come from current and former employees, or from experts in the field. [^1] There are, of course, many other possible strategies to address search frictions, such as the use of external recruiters or “raiding” the employees of competitors (see Oyer and Schaefer, 2011).
The importance of references in the labor market has been known to economists and sociologist for several decades (Rees, 1966, Granovetter, 1973). Widely cited studies put the prevalence of references as a way of filling positions well above 50%. [^2] See Ioannides and Loury (2004) and Topa (2011) for excellent reviews of the literature.
In this paper, I focus on the use of reference letters in the context of the job market for academic economists. Using an administrative data set covering all applications to a top economics department in Russia and a new prompt-based strategy to extract sentiment scores from reference letters with a large language model, I provide direct evidence that the contents of letters of recommendation are predictive of job market outcomes.
My LLM-based method has multiple advantages relative to existing methods for sentiment analysis in economics. Language models are trained to understand context, and are not focused on the presence or absence of certain words — the essence of widely applied “bag-of-words” methods. Second, by using different prompts it is possible to extract information on observable features — like the sex of the candidate — which provides a test of the model’s predictive power. Third, because language models output probabilities, the resulting sentiment scores have a clear interpretation and measurement units. Finally, prompting may be used to extract other predictive features from text, not just sentiment.
@@
The economics job market is an ideal setting to understand how references are used as a way to screen applicants. Because all job market candidates have similar formal credentials, the use of education as a signalling device is implicitly controlled for. As a way to further control for the level of training, I focus only on first time job market participants. In other words, I only look at job market candidates who are close to finishing their PhD program and have not held a permanent academic position before.
@@
The job market for economists is somewhat unique in that all open positions are widely advertised. With few exceptions, all openings are posted in a central database accessible to all market participants. Moreover, the application process itself has become highly standardized. Candidates are expected to submit a CV, a research sample (“job market paper”), and few other materials. [^3] For example, some colleges with teaching-focused positions might require a “teaching philosophy” statement and the submission of student evaluations. The extensive use of electronic applications means that applying is almost costless. In summary, the specific characteristics of the market mean that references are not being used to disseminate information about job market opportunities. [^4] For a model of the role of social networks as a manner of obtaining information about job opportunities, see Calvó-Armengol and Jackson (2004). Instead, reference letters are used as a way to screen applicants. [^5] There are alternative explanations for the use of references that do not focus on either search costs or screening. Kugler (2003) and Heath (2018) consider the possibility that firms use references as a way to lower monitoring costs and improve labor discipline. Kramarz and Skans (2014) and Wang (2013) provide evidence that in certain contexts family relations are important in finding jobs. Neither of these explanations seem particularly relevant to the economics job market.
Aside from isolating the screening channel, the economics job market is unique in that the use of references has been formalized. Indeed, job market candidates who wish to apply for a position are customarily instructed to have two or more confidential letters of reference submitted directly to the hiring committee. Because candidates typically submit the same letters to all positions they apply to, I have precise information on who wrote letters for each job market candidate in my sample. In contrast, many prior studies do not observe recommendations directly and are therefore forced to use some kind of proxy. [^6] For example, Dustmann et al. (2016) define networks in terms of ethnicity, while Bayer et al. (2008) and Hellerstein et al. (2011) use the neighborhood of residence as proxy.
@@
The issue of under what conditions indirect signaling via letters of recommendation can function efficiently has been a matter of concern in the theoretical literature. Because referees compete against each other, we can expect them to make excessive claims about the workers they are recommending in order to have them placed ahead of the candidates being recommended by a rival referee (Saloner, 1985). [^7] Other models in which referrals serve as a screening mechanism to reduce the information asymmetry inherent to the hiring process are Montgomery (1991) and Simon and Warner (1992). Moreover, referees’ motivation for referring applicants and providing truthful recommendations is not well understood. Câmara et al. (2023) model the reputation game facing referees explicitly. Remarkably, they find that referees may lie in equilibrium, and that the extent of the lying depends on the intricate interplay between preferences, reputation and lying costs. [^8] Another paper that investigates referees’ incentives theoretically is Ekinci (2016). Carmichael (1988) argues that the reason why universities have instituted a tenure system in academic departments is precisely to ensure that faculty members are properly incentivized to hire the best job market candidates.
@@
There is also some experimental evidence on the value of reference letters. Pallais (2014) conducted a field experiment on an online marketplace in which workers bid to be hired for short term employment. One of her treatment groups received detailed feedback after completing a data-entry task, while another group received only a one-to-five rating. She found that, in the two months after treatment, both treatment groups benefitted from treatment relative to a control group that was not hired. Treated workers worked more jobs and had higher earnings. In addition, making detailed positive feedback on past performance publicly available further increased their employment prospects. Abel et al. (2020) conducted an audit study where they randomly appended a reference letter to some applications submitted on behalf of job seekers to vacancies in South Africa. They found that, for the same applicant, attaching a letter increased the probability of receiving a response and the rate of interview requests by about 60 percent.
Two papers study the use of reference letters in the economics job market (Eberhardt et al., 2023, Baltrunaite et al., 2024). Their data is similar to mine in that they collect administrative hiring data from institutions hiring in the international economics job market. [^9] Eberhardt et al. (2023) study a sample of applicants to a top UK economics department, while Baltrunaite et al. (2024) look at a premier economics department and a top policy institution in Italy. While these studies do link reference letter features to job market outcomes, their focus lies more on gender differences in the linguistic contents of reference letters and their consequences.
This paper also makes a methodological contribution to the broader literature that makes use of text data in economics and finance. I show that a prompt-based strategy can take advantage of a LLM’s pre-training to extract sentiment and other scores from the text. This new strategy presents several distinct advantages over other methodologies that have been deployed in economics and finance.
In summary, I find that traditional dictionary-based methods are unable to extract information from reference letters that is predictive of job market outcomes. Prompt-based sentiment extraction, in contrast, produces sentiment scores that are strongly correlated with the probability of obtaining a permanent research position. The relationship is even stronger when predicting employment at high rank institutions.
@@
Until fairly recently, feasible strategies to take advantage of text data were rather limited. In the next section, I provide a brief review of the existing methods and a formal description of my prompt-based strategy. I then move to a description of the data, including a descriptive analysis of my prompt-based sentiment score. Section 4 presents the main results and section 5 concludes.
2 Methods: Sentiment Analysis and Prompt-based Learning
@@
In this paper, I apply a prompt-based methodology to extract sentiment and other scores from reference letters. This approach has several advantages over existing methods. In this section, I first provide a brief description of the lexical approach, which has been the most commonly applied method to assign a sentiment score to text in the economics and finance literature. I then describe the prompt-based methodology for sentiment extraction.
2.1 Dictionary Methods
Text analysis in economics has a long history. [^10] The classic Gentzkow et al. (2019) and the more recent Ash and Hansen (2023) provide great reviews of the different techniques that have been deployed to handle text in economics and finance research. In particular, several papers have studied the problem of how to assign a sentiment score to text.
@@ 2.1 Dictionary Methods
Shapiro et al. (2022) define sentiment as “a measure of the speaker’s tone, attitude, or evaluation of a topic, independent of the topic’s own sentiment orientation.” The main approach to sentiment analysis in economics and finance has been the lexical methodology, which relies on a preset dictionary of words, with each word assigned a sentiment score. [^11] Typically, words in the dictionary have a score of 1 if they are deemed to express positive sentiment, -1 if negative, and 0 for neutral words. But more complex approaches exist. Within this framework, the sentiment of a document is some simple aggregate of the scores of its constituent terms. The most common approach is to use the polarity score:
[…]
@@
where $S^{i}_{d}$ represents the sentiment score assigned by the lexicon to term $i$ of document $d$. For the simple case of ±$\pm 1$ scores, this simplifies to
@@
where $N^{+}_{d}$ and $N^{-}_{d}$ are, respectively, the counts of positive and negative terms. The methodology is often referred to as “bag of words” because the order in which the terms appear does not affect the resulting sentiment score.
@@
Another attempt at improving on the lexicon approach adds a series of heuristic rules to take into account common linguistic patterns that affect sentiment, such as negation and emphasis. A prominent example is the Vader analyzer (Hutto and Gilbert, 2014), which is designed to deal with short texts such as tweets and online reviews. In this paper, I also use SenticNet, which uses a “concept-based” lexical approach. [^12] See sentic.net for details.
@@
In addition to this limitation, bag-of-words methods also suffer from the fact that the resulting sentiment score does not have an easy interpretation or unit of measurement. While researchers customarily compare scores resulting from alternative lexicons, it is not in general possible to evaluate the relative accuracy of the different sentiment scores.
@@
A final limitation of the lexical approach is that, if the method is to be used to generate a measure other than a sentiment score, a completely new dictionary must be produced. For example, Schmader et al. (2007) created a set of dictionaries to summarize different semantic aspects of reference letters written for chemists and biochemists. Specifically, they developed sets of terms associated with the following language categories: grindstone traits, ability traits, research terms, teaching and citizenship terms, and standout adjectives. Eberhardt et al. (2023) applied a similar strategy to letters of recommendation written on behalf of job market candidates applying to a top UK economics department.
[…]

### v2

@@ Abstract
I propose a relatively simple way to deploy pre-trained large language models (LLMs) in order to extract sentiment and other useful features from text data. The method, which I refer to as prompt-based sentiment extraction, offers multiple advantages over other methods used in economics and finance. I apply my prompt-based strategy to a hand-collected corpus of confidential reference letters (RLs). I show that the sentiment contents of RLs is clearly reflected in job market outcomes. Candidates with higher average sentiment in their letters perform markedly better regardless of the measure of success chosen. Moreover, I show that disagreement among letter writers negatively affects the job market candidate’s performance. I compare my sentiment extraction approach to other commonly used methods for sentiment analysis: ‘bag-of-words’ approaches, fine-tuned language models, and querying advanced chatbots. I find that no other method can reproduce the results obtained by prompt-based sentiment extraction. Finally, I slightly modify the method to obtain ‘gendered’ sentiment scores (as in Eberhardt et al., 2023). I show that letters of reference written for female candidates emphasize ‘grindstone’ personality traits, whereas male candidates’ letters emphasize ‘standout’ traits. These gender differences negatively affect women’s job market outcomes.
JEL classification: C45; J16
Key words: Large language models; text data; sentiment analysis; reference letters.
@@ 1 Introduction
The idea that sentiment—subjective states of mind and opinions—can be an important determinant of economic decisions and outcomes is widely accepted both in micro and macroeconomic settings (e.g. Izmalkov and Yildiz, 2010, Angeletos and La’O, 2013, Angeletos et al., 2018, Maxted, 2024). It is not surprising, therefore, that considerable effort has been exerted by researchers trying to measure it. There are myriad applications of sentiment analysis in economics: in the context of business cycles (Shapiro et al., 2022), policy-level uncertainty (Baker et al., 2016, Caldara and Iacoviello, 2022), the effects of monetary policy (Ter Ellen et al., 2022), and asset pricing (García, 2013), just to name a few.
In this paper, I propose a relatively simple way to deploy pre-trained large language models (LLMs) in order to extract sentiment and other useful features from text data. I refer to this method as prompt-based sentiment extraction. In a nutshell, the method involves combining the input text with a suitable prompt, so that the LLM’s emergent abilities acquired during pre-training can be applied to the sentiment extraction task. For example, in order to recognize the emotion of the social media post “I missed the bus today!” we may append the prompt “I felt so [MASK]” and ask the LLM to fill the blank (denoted by the [MASK] token) with an emotion-bearing word. The LLM’s output, consisting of a probability distribution over its vocabulary, can then be summarized to produce the desired sentiment score.
This relatively simple method has multiple advantages over alternative techniques. Until fairly recently, the leading approaches to sentiment extraction in economics were different versions of the “bag-of-words” method, in which the order of the terms in the text do not affect the resulting sentiment score (only the presence or absence of certain words matters). In sharp contrast, within LLMs words and other language parts receive a representation that is highly dependent on the context in which the term is used. As a simple example, LLMs can distinguish without difficulty between uses of the word “fly” that refer to an insect from cases when it is used as a verb. The highly complex sequential relationships among words, clauses, sentences and paragraphs is accepted without arbitrary restrictions and in turn used to infer the correct representation of each term in the sequence.
Secondly, unlike past methods that required heavy-handed preprocessing of the text input, LLMs accept natural language “as is”, inclusive of punctuation signs and other special characters. The vocabulary size of recent LLMs is truly staggering. For example, the main model I use in my application is Llama 3.1, which has a vocabulary size of over 120 thousand tokens. For comparison, Shakespeare used around 20 thousand words in his plays and poems. In addition, Llama and other state-of-the-art LLMs are multi-lingual, accepting words in Spanish, Portuguese, Italian, German, Thai, French, and Hindi aside from English (Meta AI, 2024).
A third key advantage of prompt-based sentiment extraction is that the resulting sentiment scores have a probability interpretation. Specifically, the polarity sentiment score that I define is simply the difference between the probabilities that the LLM assigns to positive and negative completions of the prompt. This makes the interpretation of downstream applications of the scores—e.g. in regression analysis—much easier. Fourth, the LLM’s ability to correctly interpret the text can be directly assessed by testing its predictions against known text meta-data. For example, I show below that the LLM performs extremely well in predicting the job market candidate’s sex and field of research by simply “reading” the RLs. While this is not strict proof of the accuracy of the sentiment scores, it is arguably superior to the use of completely untestable alternatives. More generally, prompt-based learning can be used to extract features of text other than sentiment. For example, in this paper I use the technique to extract measures of how much the recommendation letter writers emphasize ‘grindstone’ and ‘standout’ aspects of the candidate’s personality.
I apply prompt-based sentiment extraction to a hand-collected corpus of confidential RLs submitted as a standard part of the package that academic economists submit to potential employers in academia and other sectors. Letters of recommendation have a number of desirable features that make them an ideal setting to test my sentiment extraction method. Because they are confidential, we can be sure that they have not been used as part of the training set for any of the open-source LLMs that I test-drive. As argued recently by Ludwig et al. (2025), the existence of leakage between the LLM’s training data and the economic application has the potential to severely bias results. In addition, RLs provide an interesting mix of relatively objective/factual descriptions and more opinionated content. While the structure of the letters is relatively stable, the opinions and discussion can often be quite idiosyncratic, so the task of producing a sentiment score is far from trivial. Finally, on average letters are around 2,000 words long, which is relatively easy to process for state-of-the-art LLMs.
@@
Besides showcasing prompt-based sentiment extraction, the application is meant to contribute to the growing literature on the role of references in the labor market, and how apparently innocuous small biases may contribute to gender inequality. The economics job market is an ideal setting to understand how references are used as a way to screen applicants. Because all job market candidates have similar formal credentials (a PhD), the use of education as a signalling device is implicitly controlled for. As a way to further control for the level of training, I focus only on first-time job market participants. In other words, I only look at job market candidates who are close to finishing their PhD program and have not held a permanent academic position before.
@@
The job market for economists is somewhat unique in that all open positions are widely advertised. With few exceptions, all openings are posted in a central database accessible to all market participants. The application process itself has become highly standardized. Candidates are expected to submit a CV, a research sample (“job market paper”), and at least two confidential letters of reference. [^1] Some employers may require additional materials. For example, some colleges with teaching-focused positions often require a “teaching philosophy” statement and the submission of student evaluations. Because the letter writers typically submit the same letters to all positions the job market candidates apply to (Coles et al., 2010), it is fair to assume that I have precise information on who wrote letters for each job market candidate in my sample and what the contents of those letters are. [^2] Having precise information on both the referee and the referred person is very rare. Most prior studies of referrals in the labor market do not observe recommendations directly and are therefore forced to use some kind of proxy. For example, Dustmann et al. (2016) define referral networks in terms of ethnicity, while Bayer et al. (2008) and Hellerstein et al. (2011) use the neighborhood of residence as proxy. Finally, the extensive use of electronic applications means that applying for a job opening is almost costless to the candidate. In summary, the specific characteristics of the market mean that, unlike other settings, here RLs are not being used to disseminate information about job market opportunities. [^3] For a model of the role of social networks as a manner of obtaining information about job opportunities, see Calvó-Armengol and Jackson (2004). Instead, reference letters are primarily being used as a way to screen applicants.
@@
The issue of under what conditions indirect signaling via letters of recommendation can function efficiently has been a matter of concern in the theoretical literature. Because referees compete against each other, we can expect them to make excessive claims about the candidates they are recommending in order to have them placed ahead of the candidates being recommended by rival referees (Saloner, 1985). [^4] Other models in which referrals serve as a screening mechanism to reduce the information asymmetry inherent to the hiring process are Montgomery (1991) and Simon and Warner (1992). Moreover, referees’ motivation for referring applicants and providing truthful recommendations is not well understood. Câmara et al. (2023) model the reputation game facing referees explicitly. Remarkably, they find that referees may actively lie about the candidate’s true ability in equilibrium, and that the extent of the lying depends on the intricate interplay between preferences, reputation and lying costs. [^5] Another paper that investigates referees’ incentives theoretically is Ekinci (2016). Carmichael (1988) argues that the reason why universities have instituted a tenure system in academic departments is precisely to ensure that faculty members are properly incentivized to hire the best junior job market candidates.
@@
There is also some experimental evidence on the value of reference letters. Pallais (2014) conducted a field experiment on an online marketplace in which workers bid to be hired for short term employment. One of her treatment groups received detailed feedback after completing a data-entry task, while another group received only a one-to-five rating. She found that, in the two months after treatment, both treatment groups benefitted from treatment relative to a control group that was not hired. Abel et al. (2020) conducted an audit study where they randomly appended a reference letter to some applications submitted on behalf of job seekers to vacancies in South Africa. They found that, for the same applicant, attaching a letter increased the probability of receiving a response and the rate of interview requests by about 60 percent.
A separate empirical literature has investigated whether there are differences in the language chosen to describe female and male candidates. Starting with the work of Trix and Psenka (2003) and Schmader et al. (2007), a number of papers have documented that women are described with more ‘grindstone’ attributes and fewer ‘ability’ or ‘standout’ ones. Two papers study the use of reference letters in the economics job market (Eberhardt et al., 2023, Baltrunaite et al., 2024). Their data is similar to mine in that they collect administrative hiring data from institutions hiring in the international economics job market. [^6] Eberhardt et al. (2023) study a sample of applicants to a top UK economics department, while Baltrunaite et al. (2024) look at a premier economics department and a top policy institution in Italy. They both document similar differences in language across gender lines using the lexical approach. In this paper, I confirm their findings using a prompt-based strategy.
In summary, I find that prompt-based sentiment extraction produces sentiment scores that are strongly correlated with the probability of obtaining a good job market outcome. In particular, high average sentiment among all referees are predictive of a sizeable higher probability of success. Moreover, my paper is the first to document that disagreement among letter writers can offset the positive effect of high average sentiment, which is strictly consistent with employers taking all available information in RLs pretty much at face value. While these findings are not causal, they strongly suggest that the theoretical literature might have over-emphasized the potential pitfalls of reference letters.
@@
Until fairly recently, feasible strategies to take advantage of text data were rather limited. In the next section, I provide a brief review of the existing methods and a formal description of my prompt-based strategy. I then move to a description of the data, including a descriptive analysis of prompt-based sentiment scores obtained from RLs. Section 4 presents the main results linking RL sentiment and job market outcomes. In section 5, I present results related to gender differences in the language used by letter writers to describe candidates. Section 6 concludes.
2 Sentiment Analysis and Prompt-based Learning
@@
In this paper, I present a prompt-based methodology to extract sentiment and other scores from text data. This approach has several advantages over other methods. In this section, I first provide a brief description of the lexical approach, which has been the most commonly applied method to assign a sentiment score to text in the economics and finance literature. I then briefly describe how the development of language models has revolutionized sentiment analysis and formally present the prompt-based methodology for sentiment extraction. [^7] My description of other methods is limited to the key elements necessary for a comparison with prompt-based sentiment extraction. The classic Gentzkow et al. (2019) and the more recent Ash and Hansen (2023) provide great reviews of the different techniques that have been deployed to handle text in economics and finance research. For a recent survey of “sentometrics” (the econometrics of sentiment analysis) see Algaba et al. (2020).
2.1 Lexical Approach
The basic input to any sentiment analysis task is a collection of $D$ text documents (the corpus). The raw documents must first be converted into sequences of linguistic features (tokens). The lexical approach uses word-level tokenization almost exclusively. [^8] A somewhat common exception is the use of short word sequences (2- and 3-grams) as units of analysis. Therefore, document $d$ in the corpus can be written as $(w_{d,1},\ldots,w_{d,N_{d}})$, where $N_{d}$ is the number of tokens found in the raw document.
@@ 2.1 Lexical Approach
The main approach to sentiment analysis in economics and finance has been the lexical methodology, which relies on a preset dictionary of words (a.k.a. lexicon), with each word assigned a sentiment score. [^9] Typically, words in the dictionary have a score of 1 if they are deemed to express positive sentiment, 1-1 if negative, and 0 for neutral words. But more complex approaches exist.
The main challenge with the lexical approach is creating the dictionary. Without any adjustments, the vocabulary size—the number of distinct tokens in the corpus—can be a very large number. For this reason, the lexical approach typically involves some heavy pre-processing of the text. First, all connectors, punctuation and other stopwords are eliminated from the documents. Second, words are lowercased and converted into their stem form. These steps drastically reduce the dimensionality of the problem.
Within the lexical approach, the sentiment of a document is some simple aggregate of the scores of its constituent terms. The most common approach is to use the polarity score:
$\mathrm{polarity}_{d}$ $=\frac{\sum_{i}S^{i}_{d}}{\sum_{i}\mathinner{\!\left\lvert S^{i}_{d}\right\rvert}}$,
@@
where $S^{i}_{d}$ represents the sentiment score assigned by the lexicon to term $i$ of document $d$. For the simple case of $\pm 1$ scores, this simplifies to
@@
where $N^{+}_{d}$ and $N^{-}_{d}$ are, respectively, the counts of positive and negative terms in the document. The methodology is often referred to as “bag of words” because the order in which the terms appear does not affect the resulting sentiment score. Technically, however, bag of words is a broader concept, since there are machine learning approaches to sentiment extraction that ignore word-order without using a dictionary (e.g. Frankel et al., 2022).
@@
Another attempt at improving on the lexicon approach adds a series of heuristic rules to take into account common linguistic patterns that affect sentiment, such as negation and emphasis. A prominent example is the Vader analyzer (Hutto and Gilbert, 2014), which is designed to deal with short texts such as tweets and online reviews.
@@
In addition to this limitation, bag-of-words methods also suffer from the fact that the resulting sentiment score does not have an easy interpretation or unit of measurement. Moreover, while researchers customarily compare scores resulting from alternative lexicons, it is not in general possible to evaluate the relative accuracy of the different sentiment scores.
@@
A final limitation of the lexical approach is that, if the method is to be used to generate a measure other than a sentiment score, a completely new dictionary must be produced. For example, Schmader et al. (2007) created a set of dictionaries to summarize different semantic aspects of reference letters written for chemists and biochemists. Specifically, they developed sets of terms associated with the following language categories: grindstone traits, ability traits, research terms, teaching and citizenship terms, and standout adjectives. Eberhardt et al. (2023) applied a similar strategy to letters of recommendation written on behalf of job market candidates applying to a top UK economics department. [^10] A non-trivial additional layer of complexity involves how to normalize the term counts in non-sentiment oriented dictionary tasks. Schmader et al. (2007) consider the fraction of words in a letter that belong to each of their dictionaries. Eberhardt et al. (2023) use the term-frequency-inverse-document-frequency (tf-idf) when computing the scores associated with each language category.
2.2 Fine-tuning LLMs for Sentiment Extraction
Instead of attempting to extract numerical features directly (e.g. counting tokens present in a dictionary), language models represent text tokens with dense vectors called embeddings, which are formed by having the model train on different language prediction tasks. The transformer architecture (Vaswani et al., 2017) revolutionized this approach by stacking attention layers on top of the embeddings in a way that allowed the model to learn to pick up important interactions across words in text sequences without the need to specify those dependencies explicitly. As a result, these models are able to encode the context-dependent nature of the meaning of text tokens. [^11] See Zhao et al. (2023) for a detailed review of the recent advances of LLMs.
There are two main transformer variants, with the main contrast between them being the exact way in which the attention layers are applied. Encoder-only transformers use attention layers in which every token’s embeddings are allowed to interact with every other token in the sequence. This type of model is pre-trained using masked language modeling, a prediction task in which one or more tokens in a training sequence are masked (i.e. hidden from the model). The model’s task is to predict the masked token or tokens using only the context as guidance. The great advantage of this form of training is that it is semi-unsupervised, meaning that there is no need for expensive labeled data. Instead, the model can be trained on massive corpora of raw text (e.g. Wikipedia).
Decoder-only (a.k.a. generative) transformers use “causal” attention layers, meaning that each token is only allowed to interact with the preceding tokens in the sequence. [^12] There are other transformer variants (e.g. encoder-decoders for machine translation). I focus exclusively on the different approaches to sentiment analysis. The pre-training task in this case involves predicting the correct continuation of truncated text sequences. This form of language modeling is also semi-unsupervised.
When transformers were first released, the state-of-the-art approach to using them was through a process called fine-tuning. Starting with a pre-trained language model consisting of (hundreds of) millions of trained parameters, the researcher would add to the model a simple prediction head with randomly initialized weights. The type of prediction head would depend on the exact application. Since the vast majority of parameters in the model are already trained, the researcher would only need to “fine-tune” this model. This means training the final layer for just a few epochs with a very small learning rate (other model weights are frozen during fine-tuning) on a specially collected labeled data set. The idea is to minimally modify the model to adapt it to the desired task.
To put these points formally, consider first a supervised learning NLP task that takes a text $\boldsymbol{x}$ and predicts output based on a model $\mathrm{P}\left(y\mid\boldsymbol{x};\boldsymbol{\theta}\right)$. For a text classification task, given an input text $\boldsymbol{x}$ the model predicts a label $y$ from a fixed label set $\mathcal{Y}$. For example, a sentiment analysis task might take an input such as “I loved this movie!” and predict a label $y=+$ out of the label set $\mathcal{Y}=\{+;\sim;-\}$. In this terms, the challenge of supervised learning is that in order to find optimal parameters, we need lots of supervised data $\left(\boldsymbol{x}_{i},y_{i}\right),i=1,\ldots,N$ where the $y_{i}$ are the labels (a.k.a. “ground truth”) for a number $N$ of training examples. Because language inputs are very high-dimensional [^13] Roughly, the dimensionality of the input is the vocabulary size times the length of the longest sequence the model can consider., language models require millions (or billions) of parameters to achieve good performance. The larger the number of parameters, the more supervised data is required to train the model. This is the rationale behind the use of semi-unsupervised pre-training.
It is useful to partition the parameter vector as $\boldsymbol{\theta}\equiv\left[\boldsymbol{\theta}_{pre},\boldsymbol{\theta}_{out}\right]$, where $\boldsymbol{\theta}_{pre}$ comprises the embeddings and all other model weights except the output layer. In summary, the fine-tuning approach involves using versions of these weights that were obtained through language modeling (typically by an organization with a lot of computing power) and then training only $\boldsymbol{\theta}_{out}$ through supervised learning with domain-specific data. While the approach still requires some investment in labeling cases, the investment is typically negligible compared to what would be required without transferring the pre-trained weights.
One good example of fine-tuning for sentiment analysis is the FLAIR project (Akbik et al., 2019). [^14] See also https://flairnlp.github.io/. Their model is based on distilBERT [^15] The distilBERT (Sanh et al., 2020) model is an encoder-only transformer with a maximum context size of 512 tokens. pre-trained weights. The output layer was fine-tuned using the Amazon review corpus and other similar text.
FinBERT (Huang et al., 2023) is another fine-tuned BERT-based model. Interestingly, the creators of FinBERT pre-trained the model themselves using a corpus of financial statements. They then fine-tuned it for sentiment analysis using labeled analyst reports.
The obvious limitation of fine-tuned models is their domain-specificity. Models like FLAIR or FinBERT cannot be expected to be good at predicting sentiment for documents outside the domain of their training data.